OpenStudy (anonymous):

Please help! Prove: If n is congruent to 1(mod 6), then n^2 + 2^n is composite. Please help! Prove: If n is congruent to 1(mod 6), then n^2 + 2^n is composite. @Mathematics

jimthompson5910 (jim_thompson5910):

If we're assuming that n^2 + 2^n is composite, then ab = n^2 + 2^n (mod 6) where a and b are not equal to 1 So, ab = n^2 + 2^n (mod 6) ab = 1^2 + 2^1 (mod 6) ab = 1 + 2 (mod 6) ab = 3 (mod 6) Now let a = 3 and b = 3, since ab = 3*3 = 9 = 3 (mod 6), this means that we've shown that 3 is composite mod 6 Therefore, n^2 + 2^n is composite mod 6 when n = 1 (mod 6)